The ACT Composite scores have an average of 20.8 with a standard deviation of 5.8 (this is actually true). A sample of 64 scores is selected. a. What is the probability that the sample mean will be larger than 24.1? b. What is the probability that the sample mean will be less than 16.6? c. What is the probability that the sample mean will be between 20.8 and 26.3? d. What is the probability that the sample mean will be less than 20.8? e. What is the probability that the sample mean will be larger than 31.2?

PLEASE SHOW DETAILED WORK FOR EACH PROBLEM
Answer #1

Given U 20.8 5-8 we use normal dist n=64 (> 80, large Samples) a) p (m. 24-1)= P(a-ma =P(2<0.57) 24.1- 20.8 $.8 = 0-7157. | Bd) p(x<20.8) Plzeu < 2018-2018 5.8 = P(zco) = 0.5 >31.2-20.8 e) P(x >31.27 P (au 5-8 = P(2> 1.79) = 1-P (221.79) =1-0.9633 =